# Commodification of Compute

### v1

@@ 1   Introduction
The demand for computational resources, particularly GPUs, has surged dramatically due to advancements in AI and machine learning. This increased demand has resulted in a significant supply crunch, making computational power both expensive and difficult to access for many startups and independent developers. The concept that “compute is the new oil” encapsulates this paradigm shift, highlighting the critical role that computational resources play in driving technological innovation and economic growth [1, 2].
The AI sector has experienced unprecedented growth, with substantial investments highlighting the significant economic potential and demand for computational resources. As AI adoption continues to accelerate, the need for scalable and efficient compute resource management becomes ever more critical. What is more, research indicates that the computational requirements for leading-edge AI systems, for example, have been doubling approximately every three to four months [3, 4]. Such growth rates substantially outpace Moore’s Law [5], emphasizing the urgent need for innovative solutions in resource provisioning and management [6, 7].
@@
The GCX provides a versatile platform that caters to various stakeholders in the compute market, enabling them to hedge against price volatility, secure future prices, and generate additional yield. As illustrated in Fig. (2), the GCX allows different actors to leverage its functionalities to their advantage. Alice, an AI startup founder, uses futures contracts to lock in the price of compute, safeguarding against future price increases [18]. Bob, anticipating a market correction, buys put options to secure a price floor, thereby minimizing potential losses [19]. Carol, a data center operator, sells both call and put options to collect premiums, generating additional yield. When these options expire worthless, she profits from the premiums, smoothing her revenue stream (producer hedge) [20]. By utilizing the GCX, these actors can effectively manage their risks and optimize their financial strategies in the dynamic compute market.
In all cases, notice that the GCX enables various expressions of market sentiment, allowing participants to profit and protect against the evolution of compute prices, whether they anticipate an increase, decrease, or sideways movement. This flexibility ensures that the GCX can benefit a diverse range of market participants, each with their unique perspectives and strategies.
@@ 3   Commodification of Compute
The core value proposition of CaaC boils down to ready availability of affordable compute capacity for organizations on tap, much like any utility service for modern life such as electricity or piped gas. Computing transitions from an owned asset requiring upfront capital allocation and skilled labor, to an instantly usable operational expenditure that auto-scales to workload needs dynamically [25].
@@
Thirdly, it provides flexibility and efficiency. Users can purchase compute power as needed, allowing for flexible scaling of resources based on demand. This dynamic allocation helps optimize resource usage and reduce wastage [36, 41]. Moreover, access to a marketplace for compute resources means that businesses and individuals can acquire the exact amount of compute power they need, when they need it, without long-term commitments [42, 43].
Lastly, commodification promotes accessibility and inclusivity. Compute commodification democratizes access to high-performance computing, allowing smaller businesses and individuals to leverage powerful compute resources that were previously accessible only to large corporations [35, 44]. Easier access to compute power can spur innovation and development, enabling new applications and technologies that rely on intensive computational resources [45, 3].
@@ 6   Learning from Power to Enhance Compute Resource Trading
Table 1 provides a side-by-side comparison of key characteristics between power and compute markets.
@@ 8.1.3 Clearing Layer
The Clearing Layer facilitates smooth and secure trade settlements. It includes guarantors who provide trade guarantees and manage the delivery of compute resources. Guarantors also serve as gatekeepers to the GCX, ensuring market participants maintain sufficient collateral to cover their risk both pre- and post-trade. Key components include:
@@ 8.2   Process Flow: Clearing and Guarantee Pools
Figure 5: Workflow of the GCX App for Short and Long Futures: This process starts with the expiration of a future contract, either short in the GCX App or long in a company-built app. Guarantors match participants to deliver the compute resources. The short participant is required to deliver to the long. If the short can deliver, the buyer pays, and the compute is delivered. If the short fails to deliver, their staked tokens are slashed, and the buyer is compensated. The reputation score and risk engine are updated in the risk management layer accordingly.

### v2

@@ 1   Introduction
The demand for computational resources, particularly GPUs, has surged dramatically due to advancements in AI and machine learning, see Tab. (1) and Fig. (2) [1]. This increased demand has resulted in a significant supply crunch, making computational power both expensive and difficult to access for many startups and independent developers. From a16z in [2]: “There is no sign that the GPU shortage we have today will abate in the near future.” Nvidia argues that the shortage of compute resources is not only due to a lack of GPUs themselves but rather their unavailability due to being tied up in contracts [3].
The ongoing compute shortages have compelled companies to innovate in order to maintain access to essential resources with firms seeking to reduce their GPU requirements. In 2023 alone, investors channeled hundreds of millions of dollars into startups focused on developing software that maximizes the efficiency of existing GPU resources. According to Chetan Kapoor from AWS [4], “If there’s an ask from a particular customer that needs 1,000 GPUs tomorrow, that’s going to take some time for us to slot them in. But if they are flexible, we can work it out.” This flexibility and innovation are crucial as the industry navigates the challenges posed by limited GPU availability [5].
The rapid growth of ChatGPT [6] and the AI sector has significantly stressed the semiconductor industry and its supply chain. Companies that anticipated this demand reserved GPUs early, causing prices to double since 2020. Consequently, while spot instance pricing is available, vendors struggle to fulfill allocations, and acquiring clusters is nearly impossible without insider connections. This surge in software demand has outpaced the physical infrastructure’s ability to produce hardware. The increasing complexity of chips, high-performance networking, and packaging has escalated prices, increased failure rates, and reduced yields. Nvidia’s [7] Robert Ober notes that despite cloud leaders’ requests to boost production tenfold, the intricate nature of these systems prevents such rapid scaling. Improving maximum performance through optimized “model flop utilization” and securely intermixing users to ensure constant hardware operation is essential as we work towards scaling up compute manufacturing [3].
The concept that “compute is the new oil” encapsulates this paradigm shift, highlighting the critical role that computational resources play in driving technological innovation and economic growth [8, 9].
Figure 2: Trends in FLOPs for 121 milestone ML models between 1952 and 2022. The graph illustrates the growth in computational requirements (FLOPs) over time, highlighting three distinct eras. A noticeable change in slope occurs around 2010, corresponding to the advent of Deep Learning. Additionally, a new large-scale trend emerges in late 2015, indicating further advancements in ML model complexity. Replicated from [1].
Figure 1: Compute trends across three eras of machine learning. The table presents the periods, the data set size, the scale of compute from start to end, and the corresponding doubling times. In the Pre Deep Learning [10] era (1952 to 2010), all models exhibit a scale increase from 0.00003 to 200 TFLOPs with a doubling time of 21.3 months. The Deep Learning era (2010 to 2022) saw regular-scale models increase from 700 to 2,000,000 TFLOPs with a significantly faster doubling time of 5.7 months. Finally, the Large-Scale era (September 2015 to 2022) involved large-scale models scaling from 4,000,000,000 to 800,000,000,000 TFLOPs with a doubling time of 9.9 months. From [1].
As AI, and general cloud compute, adoption continues to accelerate, the need for scalable and efficient compute resource management becomes ever more critical. What is more, research indicates that the computational requirements for leading-edge AI systems, for example, have been doubling approximately every three to four months [11, 12], see again Tab. (1). Such growth rates substantially outpace Moore’s Law [13], emphasizing the urgent need for innovative solutions in resource provisioning and management [14, 15].
@@
The GCX provides a versatile platform that caters to various stakeholders in the compute market, enabling them to hedge against price volatility, secure future prices, and generate additional yield. As illustrated in Fig. (2), the GCX allows different actors to leverage its functionalities to their advantage. Alice, an AI startup founder, uses futures contracts to lock in the price of compute, safeguarding against future price increases [26]. We already see glimpses of this happening albeit in a fragmented and inefficient way compared to having a centralized marketplace: In 2023, for example, Resemble turned to FluidStack, a tiny provider that welcomes one-week or one-month GPU reservations, and has recently joined San Francisco Compute Group, a consortium of startups jointly committing to buy and split GPU capacity [5]. Bob, anticipating a market correction, buys put options to secure a price floor, thereby minimizing potential losses [27]. Carol, a data center operator, sells both call and put options to collect premiums, generating additional yield. When these options expire worthless, she profits from the premiums, smoothing her revenue stream (producer hedge) [28]. By utilizing the GCX, these actors can effectively manage their risks and optimize their financial strategies in the dynamic compute market.
Unlike commodities such as oil and coffee, compute has not yet become a tradable commodity in established markets. The absence of compute commodity markets may be attributed to several factors, including the rapid evolution of technology, the complexity and diversity of compute resources, and the lack of standardized units for trading compute. In all cases, notice that the GCX enables various expressions of market sentiment, allowing participants to profit and protect against the evolution of compute prices, whether they anticipate an increase, decrease, or sideways movement. This flexibility ensures that the GCX can benefit a diverse range of market participants, each with their unique perspectives and strategies.
@@ 3   Commodification of Compute
The core value proposition of CaaC boils down to ready availability of affordable compute capacity for organizations on tap, much like any utility service for modern life such as electricity or piped gas. Unlike traditional cloud providers such as AWS, which require significant setup and management, CaaC presents compute resources as an instantly usable operational expenditure. This model dynamically scales to workload needs without extensive configuration, reducing both the technical expertise required and the administrative burden on users. Computing transitions from an owned asset requiring upfront capital allocation and skilled labor, to an instantly usable operational expenditure that auto-scales to workload needs dynamically [33].
@@
By activating dormant resources, the compute market could significantly expand. There is substantial underutilized computational power in personal devices like smartphones and PCs, which could be leveraged through financial incentives. Similarly, many companies have in-house compute resources that are often idle during non-peak hours, representing another untapped resource. Additionally, the commodification model could lower barriers for smaller providers, encouraging the emergence of startups and niche providers. This would foster a more competitive and diverse market, attracting investment and driving innovation.
To estimate the potential market size, consider the following contributions:
  • Personal Devices: With an estimated 6.84 billion smartphones globally, contributing 1 GFLOP for 2 hours daily, the total annual revenue could reach \$499.32 billion based on current rates for cloud services.
  • Personal Computers: With an estimated 2 billion personal computers globally, contributing 10 GFLOP for 2 hours daily, the total annual revenue could reach \$1.46 trillion.
  • In-House Company Resources: If multiple corporations collectively contribute 1 exaFLOP of their compute capacity continuously, the annual revenue could be \$87.6 billion.
  • Emerging Smaller Providers: Assuming 10,000 new providers each contribute 100 TFLOPs, the annual revenue could also be \$87.6 billion.
Summing these contributions, the potential market size could reach more than \$2 trillion annually. This substantial economic potential underscores another significant benefit of commodifying compute resources.
Thirdly, commodification also provides flexibility and efficiency. Users can purchase compute power as needed, allowing for flexible scaling of resources based on demand. This dynamic allocation helps optimize resource usage and reduce wastage [44, 49]. It is estimated that about 76% of non-production cloud resources (such as those used for development, staging, testing, and QA) remain idle during non-working hours. This translates to a substantial amount of compute resources being paid for without being actively used. Also, many instances are provisioned with more capacity than necessary. For example, around 40% of cloud instances are sized larger than needed, leading to unnecessary expenses. Downsizing these instances could save significant costs [50, 51]. Access to a marketplace for compute resources means that businesses and individuals can acquire the exact amount of compute power they need, when they need it, without long-term commitments [52, 53].
Lastly, commodification promotes accessibility and inclusivity. Compute commodification democratizes access to high-performance computing, allowing smaller businesses and individuals to leverage powerful compute resources that were previously accessible only to large corporations [43, 54]. Easier access to compute power can spur innovation and development, enabling new applications and technologies that rely on intensive computational resources [55, 11]. Quoting from [5]: “The startup ecosystem is trying to get together and try to figure out ‘How do we battle, how do we fight for compute?’ Otherwise, it would be a really unfair game. Prices are just too high.”
@@ 6   Learning from Power to Enhance Compute Resource Trading
We note also that, in essence, the drive towards the commodification of computing resources is primarily motivated by the need to enhance insights into future cash flows and mitigate the risks associated with uncertain future utilization levels. It’s less about the actual computing power and more about providing financial predictability and stability. Data centers are becoming the new power plants in this evolving landscape [79].
@@
Tab. (2) provides a side-by-side comparison of key characteristics between power and compute markets.
@@ 7   Standardizing Compute Units
Our goal is to commodify compute resources by creating a universal metric that can be applied across a wide range of computational tasks. This doesn’t mean that one reference system applies to all tasks uniformly. Instead, it means that we have a standard approach to measure and compare computational work by normalizing against task-specific reference systems. In essence, the Compute Hours metric is intended to provide a standardized measure of computational work that can be applied broadly, while still respecting the unique characteristics of different types of tasks.
@@ 8.1.3 Clearing Layer
The Clearing Layer facilitates smooth and secure trade settlements. It includes guarantors who provide trade guarantees and manage the delivery of compute resources. Guarantors also serve as gatekeepers to the GCX, ensuring market participants maintain sufficient collateral to cover their risk both pre- and post-trade.
To address potential concerns regarding the requirement for compute providers to post collateral, it is important to highlight the significant benefits of this approach. Requiring collateral helps mitigate counterparty risk, ensuring that market participants fulfill their obligations, thereby enhancing market stability and integrity. This requirement increases buyer confidence, fostering trust and encouraging market participation. While posting collateral introduces an upfront cost for compute providers, it standardizes risk management practices, leading to more efficient operations. Additionally, collateral enables larger and more complex transactions, promoting market growth and higher profitability for compute providers. Though the collateral requirement is a downside, it enables the market to function effectively, resulting in a net positive for participants. Therefore, the collateral requirement strikes a balance between ensuring market integrity and providing a secure platform for compute providers, ultimately leading to long-term benefits that outweigh the initial costs. Key components include:
@@ 8.2   Process Flow: Clearing and Guarantee Pools
Figure 6: Workflow of the GCX App—Fig. (1)—for Short and Long Futures: This process starts with the expiration of a future contract, either short in the GCX App or long in a company-built app. Guarantors match participants to deliver the compute resources. The short participant is required to deliver to the long. If the short can deliver, the buyer pays, and the compute is delivered. If the short fails to deliver, their staked tokens are slashed, and the buyer is compensated. The reputation score and risk engine are updated in the risk management layer accordingly.